A project manager is working on a large construction project. His plan says that the project should end up costing $1.5 million, but he’s concerned that he’s not going to come in under budget. He’s spent $950,000 of the budget so far, and he calculates that he’s 57% done with the work, and he doesn’t think he can improve his CPI above 1.05. What is the TCPI? Furthermore, combined with the earlier TCPI, what does this say about the project?
Answer #1

The Budgeted actual cost(BAC) = $1500000

The percentage complete = 57%

The earned value (EV)= 0.57*1500000=$855000

Actual cost(AC)= $950000

TCPI= (BAC-EV)/(BAC-AC)=(1500000-855000)/(1500000-950000)= 1.17

This means that the To-complete performance index (TCPI) of 1.17 is needed to take the project under budget but as the manager calculated that the CPI is 1.05 the project will cross above the budget.

Please comment if any doubt

Rate me

Thanks